If 1.25 g of sodium carbonate was dissolved in water and the solution titrated with 22.7 ml of hydrochloric acid to a medal orange end point, calculate (i) the molarity and (ii) the percent by mass (density of sution = 1.02 g/ml) of the hydrochloric acid solution

1
Expert's answer
2022-05-18T13:52:07-0400

Na2CO3 + 2 HCl => 2 NaCl + CO2 + H2O

1. According to the above reaction, 1 mole (106 g) of Na2CO3 reacts with 2 moles of HCl. we have 1.25 g of Na2CO3 and we can find how many moles of HCl react with it:

106 gr —> 2 mol

1.25 gr —> x mol

x = 1.25 * 2 / 106 = 0.0236 mol.

2. We can now find the molar concentration.To do this, divide the amount of solute (HCl) by the volume of solution (L):

CM = n / V = 0.0236 mol / 0.0227 L ≈ 1.04 mol/l.

3. We can now find the percentage concentration of the solution. To do this, we need to divide the mass of solute (HCl) by the mass of the solution and multiply by 100:

Percent by mass= Mass of Solute/ Mass of Solution * 100


Massa of Solution = V (volume) * p (density)=22.7 * 1.02 = 23.154 gr .

Percent by mass = 0.86 gr / 23.154 gr * 100= 3.71 %

ANSWER: Molarity = 1.04 mol/l

Percent by mass = 3.71% .